pH

I have a beaker with 250.0 mL of distilled water.  I dissolve 10.0 g of sodium nitrite (NaNO2) and 10.0 g of barium nitrite [Ba(NO2)2].  What is the pH of the resulting solution?
Answer #1

pH=pKa+log(b/a)

pH=4.77+ log(10/68.99)/(10/229.35) = 4.31

source: instructor answer key
answered by: Tommy Cain